OpenStudy (anonymous):

Solve it with the Substitution method 3x+5y=-1

OpenStudy (anonymous):

You cannot solve this system by substitution with simply one equation. You are missing one.

OpenStudy (anonymous):

Ok so first you can take y=2x-8 and substitute it into 3x+5y=-1. Now it looks like this: 3x +5(2x-8)=-1. Now distribute. 3x+10x-40 = -1. Add like terms. 13x-40=-1. Add 40 to both sides. 13x = 39. Divide both sides by 13. x = 13. So now we have the x coordinate (13, y) but we still need the Y. As one equation stated, y=2x-8. So y=2(13)-8. Distribute. y=26-8 And subtract. y=18. The solution is (13, 18).

OpenStudy (anonymous):

lucasmedina10, it won't satisfy the given equations.

OpenStudy (anonymous):

True, pat18. Maybe the given equations were written incorrectly?

OpenStudy (anonymous):

yeah, i had another solution but it can't equate the equations.